# Posterior SBC: Simulation-Based Calibration Checking Conditional on Data

### v1

@@ 3.1 Simple hierarchical model – choice of model parameterization
We implement this model with centered and non-centered parameterizations using the Stan probabilistic programming language (Stan Development Team,, 2023) and run the posterior inference with its default MCMC sampling algorithm, the no-U-turn sampler (Hoffman and Gelman,, 2014; Stan Development Team,, 2023). Despite the slower sampling, we raise the target acceptance ratio, called adapt delta, from $0.80$ to $0.99$, to more reliably sample posteriors with highly varying curvature. In each iteration of both prior and posterior SBC, we sample four MCMC chains of 1000 warm-up draws and 1000 posterior draws.
@@ Results
Figure 3: Prior SBC checking for the hierarchical model with (a) centered and (b) non-centered parameterization using four different test quantities. We use PIT-ECDF difference plots with 95% simultaneous confidence intervals under the assumption of calibration. Neither parameterization shows calibration issues.
@@
Figure 4: Posterior SBC for the hierarchical model with an observed data with large $\tau$ and small $\sigma$, corresponding to weak population prior and strong likelihood, implying a funnel shaped posterior with the non-centered parameterization. The PIT-ECDF plots for the non-centered parametrization show that the right tail of the posterior for $\mu$ tends to be too thin. The inference for $\tau$ shows a similar issue.
@@
Figure 5: Posterior SBC for the hierarchical model with an observed data with small $\tau$ and large $\sigma$, corresponding to strong population prior and weak likelihood, implying a funnel shaped posterior with the centered parametrization. The PIT-ECDF plots for the centered parameterization show that the inference often overestimates $\tau$.
@@
To assess the calibration of the inference, we inspect the uniformity of the PIT values of the joint log-likelihood, and of the parameters $(\mu_{0}$, $\tau$, $\sigma)$ shared between the groups. We employ the graphical uniformity test introduced by Säilynoja et al., (2022), see e.g., Figure 3. The test shows 95% simultaneous central confidence bands for the empirical cumulative distribution function (ECDF) of the discrete PIT values under the assumption of calibration. The PIT values for a given quantity are obtained by transforming the quantity with respect to the corresponding posterior draws, see Figure 2. That is, for the population parameters, the prior draw is transformed with respect to the posterior draws. For the joint log-likelihood, we compare the log-likelihood of the sample, when evaluated with the known true parameter values, against the joint log-likelihood evaluated with the parameter posterior draws.
@@
For prior SBC, Figure 3 shows no noticeable calibration issues in the inference with either parameterization. This would indicate to the modeller that everything is fine with either of these model implementations.
@@
Next, we assess the calibration via posterior SBC by conditioning on the dataset with large $\tau$ and small $\sigma$. This corresponds to the weak population prior and strong likelihood case, implying a funnel shaped posterior with the non-centered parameterization. Running 500 iterations of posterior SBC reveals a very different story for the calibration of the two parameterizations. As shown in Figure 4, the centered parameterization shows excellent calibration, but the non-centered parameterization has calibration issues when evaluating the population level parameters $\mu_{0}$ and $\tau$
@@
With the second dataset, generated with small $\tau$ and large $\sigma$ corresponding to a case of strong population prior and weak likelihood, implying funnel shape for the posterior of the parameters in the model with centered parameterization. The posterior exhibits highly varying curvature, making it hard to explore for the inference algorithm. When we run posterior SBC conditional on this data, the presence of calibration issues with the centered parameterization of the model is confirmed. Figure 5 shows the calibration assessment of posterior SBC, revealing the clear calibration issues with the centered parameterization, but also showing possible overestimation of $\tau$ in the inference run with the non-centered parameterization. If we would like to inspect this effect more closely, we could run posterior SBC with more iterations to increase the sensitivity of the calibration checking.
@@
In the presented cases, Stan interface conducted automated inference diagnostics, producing warnings for one or both of the parameterizations. In prior SBC both of the model parameterizations faced a considerable number of iterations with high $\hat{R}$ values (Vehtari et al.,, 2021), implying bad mixing of the MCMC chains. Additionally, $12\%$ of the iterations of the centered parameterization model had one or more divergent transitions, implying possibly biased posterior inference. In posterior SBC, the first dataset with strong likelihood produced warnings for the non-centered parameterization, and the second dataset with weak likelihood was problematic for the centered parameterization. These warnings align with our findings with SBC by also indicating the inference problems. This is plausible since the issues are caused by the challenging posterior geometry. However, good diagnostics are not necessarily available for new inference methods, or as shown in Section 3.3, when the model complexity grows, the inference issues might be more subtle, but still have a strong impact on the posterior estimates and calibration.
@@ 3.2 Lotka-Volterra model – focusing computational efforts
where $\hat{H}_{t}$ and $\hat{L}_{t}$ are the recorded numbers of pelts collected, $H_{t}$ and $L_{t}$ the latent true (trappable) populations of the species at that time, and $\sigma_{h}$ and $\sigma_{l}$ the error terms related to the species. All of these parameters are constrained to be positive. We don’t have strong intuition on the interactions between the Lotka-Volterra model parameters, so we assign them the following independent priors, used for example in the case study of Carpenter, (2018),
@@
Figure 6: Prior predictive draws of the observed pelt counts over time. Top left shows the true historical observation data. Many prior draws don’t exhibit the observed periodicity, or are in other ways unrealistic.
@@ Results
In prior SBC, we observe some prior predictive draws resulting in posterior inference with severe convergence issues. When investigating the calibration of the inference, we look at the PIT-ECDF of the joint log-likelihood in Figure 8. There, we see some potential calibration issues manifesting as fewer than expected PIT values between 0.5 and 1. The calibration assessment for the individual parameters can be found in the appendix, but shows no calibration issues for the individual parameters before corrections for multiple testing.
Posterior SBC, in turn, indicates good calibration for the observed historical data. A graphical assessment of the uniformity of the PIT joint log-likelihood values is shown in Figure 8. The faster inference during the posterior SBC allowed us to iterate on the model and experiment on using Pathfinder for initializing the MCMC chains in the more desirable posterior mode. More details are shown in Appendix A.1.
@@
Figure 8: PIT-ECDF plots of the posterior joint log-likelihood in the Lotka-Volterra model for both prior, and posterior SBC. Prior SBC uses only 250 iteration, resulting in wider confidence bands. Posterior SBC could include 500 iterations due to faster computation.
@@ 3.3 Posterior SBC in amortized Bayesian inference
In this case study, we show how posterior SBC can serve as a powerful diagnostic for amortized Bayesian inference workflows. Amortized Bayesian inference (see Zammit-Mangion et al.,, 2024, for an overview) constitutes a new wave of Bayesian inference methods, where neural networks learn a direct surrogate for the posterior distribution. Normalizing flows (Papamakarios et al.,, 2021) are arguably the most prominent neural density estimator for amortized inference, but other architectures have recently been explored as well, such as score-based diffusion models (Geffner et al.,, 2023; Sharrock et al.,, 2024), flow matching (Dax et al.,, 2023), and consistency models (Schmitt et al., 2023b, ).
@@
Amortized inference consists of two stages: (1) A training stage, where a generative neural network $q_{\phi}$ learns to distill relevant information from the probabilistic model based on synthetic data $(\theta,y)\sim p(\theta,y)$ from the joint model; and (2) the inference stage, where the trained neural networks approximate the posterior distribution for an arbitrary new dataset $y_{\mathrm{obs}}$. Amortized inference casts fitting a probabilistic model as a neural network prediction task, which typically takes well under a second. This achieves near-instant approximate posterior sampling $\theta\sim q_{\phi}(\theta\,|\,y_{\mathrm{obs}})$ and even direct posterior density evaluations for arbitrary parameter values $\theta$.
@@
where $\varepsilon\sim\mathcal{N}(0,1)$ is Gaussian noise and $\Delta t$ controls the time resolution (5 milliseconds in our data model). The DDM is additionally parameterized by a decision threshold $\alpha$, an evidence starting point $\beta$ to represent biases, and a non-decision time $\tau$ to account for motor reaction times. We refer to Appendix A.2 for more details.
In a line of applied work, neurological research has identified neural markers that are associated with decision-making processes. In a recent paper, Ghaderi-Kangavari et al., (2023) propose a set of multiple statistical models that integrate both the cognitive drift-diffusion model and a compressed representation of neural measurements, as found in EEG or fMRI data. Such integrative models, which combine neural processes at the level of single experimental trials, represent the current state-of-the-art in cognitive modeling research. This case study implements two probabilistic models $M_{1}$ and $M_{2}$, which correspond tomodels #2 and #6 from Ghaderi-Kangavari et al., (2023). We use deep neural networks to fit an amortized posterior approximator for each model. Detailed specifications of models and neural networks can be found in Appendix A.2. We showcase how posterior SBC can yield important information to compare the neural approximations of the candidate models conditional on two real datasets from another study (Georgie et al.,, 2018).
@@ Results
In the following, we summarize the main results of the case study and refer to Appendix A for additional details. In neural amortized inference, we must check the inference validity for different numbers of observations because the latter is passed as a conditioning variable to the inference network $f_{\phi}$. To this end, we confirm that amortized inference on synthetic datasets simulated from the joint model $p(N,\theta,y)$ can recover the ground-truth values of the model parameters that are identifiable for both $N_{\text{obs}}$ and $2N_{\text{obs}}$ (see Appendix A). Further, we can confirm that prior SBC checking for datasets with $N_{\text{obs}}$ as well as $2N_{\text{obs}}$ yields satisfactory results (see Figure 10). However, this conclusion drastically changes when we perform SBC checking conditional on the real datasets $y_{\text{obs}}^{(1)}$ and $y_{\text{obs}}^{(2)}$. As shown in 10(a), posterior SBC indicates minor data-conditional calibration issues of model $M_{1}$ for the first observed dataset $y_{\text{obs}}^{(1)}$ while calibration conditional on the second observed dataset $y_{\text{obs}}^{(2)}$ is within the acceptance band. The second model $M_{2}$, however, shows pathologically bad calibration of the mean visual encoding latency $\mu_{(e)}$ conditional on either observed dataset (see 10(b)). As a consequence, model $M_{2}$ might be refined according to the iterative Bayesian workflow (Gelman et al.,, 2020; Schad et al.,, 2020) with a particular focus on the mean visual encoding latency parameter $\mu_{(e)}$. In this case study, posterior SBC could flag calibration issues of an amortized approximator that were not detectable with prior SBC, which illustrates its potential as a default data-conditional diagnostic in amortized Bayesian workflows (Radev et al.,, 2023; Schmitt et al.,, 2024).
@@ 4 Discussion
Prior SBC and posterior SBC have different roles in the Bayesian workflow. Posterior SBC can be used to diagnose whether the used inference algorithm is sampling faithfully from the posterior and the augmented posteriors, which are close to the original posterior. If there is a worry that the augmented posterior is too different from the original posterior having double data size, for it is possible to use part of the data for the first posterior and make the augmented data size to be more similar with the original data size.
Passing posterior SBC checking is a necessary but not sufficient condition to trust the posterior inference in the same way as most convergence diagnostics given finite computation time. If the common MCMC diagnostics like divergences and $\widehat{R}$ indicate big problems, it is good to first investigate the possible reasons before spending computation time for posterior SBC.

### v2

@@ 4.1 Simple hierarchical model – choice of model parameterization
We implement this model with centered and non-centered parameterizations using the Stan probabilistic programming language (Stan Development Team,, 2023) and run the posterior inference with its default MCMC sampling algorithm, the no-U-turn sampler (Hoffman and Gelman,, 2014; Stan Development Team,, 2023). Despite the slower sampling, we raise the target acceptance ratio, called adapt delta, from $0.80$ to $0.99$, to more reliably sample posteriors with highly varying curvature. In each iteration of both prior and posterior SBC, we sample four MCMC chains of 1000 warm-up draws and 1000 posterior draws. To automate SBC iterations, we used the sbc R package (Kim et al.,, 2024).
@@ Results
Figure 3: Prior SBC checking for the hierarchical model with (a) the centered and (b) the non-centered parameterization using four different test quantities. Subplots show PIT-ECDF difference plots using four different test quantities and 95% simultaneous confidence intervals under the assumption of uniformity. As all blue PIT-ECDF difference lines are inside the 95% simultaneous confidence intervals, the inference seems to be calibrated when SBC iterations are averaged over prior draws.
@@
Figure 4: Posterior SBC for the hierarchical model with an observed data with large $\tau$ and small $\sigma$, corresponding to weak population prior and strong likelihood, implying a close to normal posterior with (a) the centered parameterization, and a funnel shaped posterior with (b) the non-centered parameterization. Subplots show PIT-ECDF difference plots using four different test quantities and 95% simultaneous confidence intervals under the assumption of uniformity. The blue PIT-ECDF difference lines for the centered parameterization stay inside the envelope, indicating calibrated inference. The blue PIT-ECDF difference lines for the non-centered parametrization show that the right tail of the approximated posterior for $\mu_{0}$ and $\tau$ tends to be thin (the line dips outside of the envelope) when SBC iterations are averaged over the observed data posterior draws.
@@
Figure 5: Posterior SBC for the hierarchical model with an observed data with small $\tau$ and large $\sigma$, corresponding to strong population prior and weak likelihood, implying a funnel shaped posterior with (a) the centered parametrization, and a close to normal posterior with (b) the non-centered parametrization. Subplots show PIT-ECDF difference plots using four different test quantities and 95% simultaneous confidence intervals under the assumption of uniformity. The blue PIT-ECDF difference lines for the centered parameterization show that the inference often overestimates $\tau$ (the line dips below the envelope), when SBC iterations are averaged over the observed data posterior draws. The blue PIT-ECDF difference lines for the non-centered parameterization stay inside the envelope, indicating calibrated inference.
@@
To assess the calibration of the inference, we inspect the uniformity of the PIT values of the joint log-likelihood, and of the parameters $(\mu_{0}$, $\tau$, $\sigma)$ shared between the groups. PIT values are computed as described in Sections 2 and 3. We employ the graphical uniformity test introduced by Säilynoja et al., (2022), see e.g., Figure 3. The test compares the empirical cumulative distribution function (ECDF) of the discrete PIT values to 95% simultaneous central confidence bands under the assumption of discrete uniformity. We use the implementation of the test available in the bayesplot R package (Gabry and Mahr,, 2024). To provide better dynamic range in the plots, we show the ECDF difference (the observed ECDF minus the expected CDF values for uniform distribution) as recommended by Säilynoja et al., (2022). The PIT values for a given quantity are obtained by transforming the quantity with respect to the corresponding posterior draws, see Figure 2. That is, for the population parameters, the prior draw is transformed with respect to the posterior draws. For the joint log-likelihood, we compare the log-likelihood of the sample, when evaluated with the known true parameter values, against the joint log-likelihood evaluated with the parameter posterior draws.
@@
After prior SBC with 500 iterations, Figure 3 shows no noticeable calibration issues in the inference with either parameterization. This would indicate to the modeller that everything is fine with either of these model implementations.
@@
Next, we assess the calibration via posterior SBC by conditioning on the dataset with large $\tau$ and small $\sigma$. This corresponds to the weak population prior and strong likelihood case, implying a funnel shaped posterior with the non-centered parameterization. Running 500 iterations of posterior SBC reveals a very different story for the calibration of the two parameterizations. As shown in Figure 4, the centered parameterization shows excellent calibration, but the non-centered parameterization has calibration issues when evaluating the population level parameters $\mu_{0}$ and $\tau$.
@@
The second dataset has been generated with small $\tau$ and large $\sigma$. This corresponds to a case of strong population prior and weak likelihood, implying a funnel shape for the posterior of the parameters in the model with centered parameterization. The posterior exhibits highly varying curvature, making it hard to explore for the inference algorithm. When we run posterior SBC conditional on this data, the presence of calibration issues with the centered parameterization of the model is confirmed. Figure 5 shows the calibration assessment of posterior SBC, revealing the clear calibration issues with the centered parameterization, but also showing possible overestimation of $\tau$ in the inference run with the non-centered parameterization. If we would like to inspect this effect more closely, we could run posterior SBC with more iterations to increase the sensitivity of the calibration checking.
@@
In the presented cases, the Stan interface conducted automated inference diagnostics (partially in Stan itself and partially using the posterior R package, Bürkner et al.,, 2024), producing warnings for one or both of the parameterizations. In prior SBC both of the model parameterizations faced a considerable number of iterations with high $\widehat{R}$ values (Vehtari et al., 2021b, ), implying bad mixing of the MCMC chains. Additionally, $12\%$ of the iterations of the centered parameterization model had one or more divergent transitions, implying possibly biased posterior inference. In posterior SBC, the first dataset with strong likelihood produced warnings for the non-centered parameterization, and the second dataset with weak likelihood was problematic for the centered parameterization. These warnings align with our findings with SBC by also indicating the inference problems. This is plausible since the issues are caused by the challenging posterior geometry. Although in this specific case the convergence diagnostics did indicate issues, the $\widehat{R}$ and divergence diagnostics do not quantify the direction or magnitude of bias for each parameter, which on the other hand can easily be read from the graphical PIT plots. For example, Figure 5 shows that in the centered parameterization case the inference tend to overestimate $\tau$ (dip in the left hand side of PIT ECDF-difference plot indicates missing posterior mass for small values of $\tau$), but the inference for $\sigma$ is reasonably calibrated. Furthermore, good diagnostics are not necessarily available for new inference methods, which is demonstrated with amortized Bayesian inference in Section 4.3. Finally, when the model complexity grows, the inference issues might be more subtle, but still have a strong impact on the posterior estimates and calibration.
@@ 4.2 Lotka-Volterra model – focusing computational efforts
where $\hat{H}_{t}$ and $\hat{L}_{t}$ are the recorded numbers of pelts collected, $H_{t}$ and $L_{t}$ the latent true (trappable) populations of the species at that time, and $\sigma_{h}$ and $\sigma_{l}$ the error terms related to the species. All of these parameters are constrained to be positive. We do not have strong intuition on the interactions between the Lotka-Volterra model parameters, so we assign them the following independent priors, used for example in the case study of Carpenter, (2018),
@@
Figure 6: Prior predictive draws of the observed pelt counts over time. Top left shows the true historical observation data. Many prior draws do not exhibit the observed periodicity, or are in other ways unrealistic.
@@ Results
In prior SBC, we observe some prior predictive draws resulting in posterior inference with severe convergence issues. When investigating the calibration of the inference, we look at the PIT-ECDF of the joint log-likelihood in Figure 8. There, we see some potential calibration issues manifesting as fewer than expected PIT values between 0.5 and 1. The calibration assessment for the individual parameters can be found in the supplementary material, but shows no calibration issues for the individual parameters before corrections for multiple testing.
Posterior SBC, in turn, indicates good calibration for the observed historical data. A graphical assessment of the uniformity of the PIT joint log-likelihood values is shown in Figure 8. The faster inference during the posterior SBC allowed us to iterate on the model and experiment on using Pathfinder for initializing the MCMC chains in the more desirable posterior mode. More details are shown in the supplementary material at Appendix A.
@@
Figure 8: PIT-ECDF plots of the posterior joint log-likelihood test quantity in the Lotka-Volterra model for both prior and posterior SBC. Prior SBC uses only 250 iterations, resulting in wider confidence bands. Posterior SBC could include 500 iterations due to faster computation. The blue PIT-ECDF difference line for the prior SBC dips below the envelope, indicating the inference is not well calibrated when averaged over the prior draws. The blue PIT-ECDF difference line for the posterior SBC stays inside the envelope, indicating the inference is well calibrated when averaged over the observed data posterior draws.
O’Brien et al., (2025) demonstrate a case where numerical error in the ODE solver distorts the posterior. From the usual inference diagnostic perspective, the bias due to the distortion is not distinguishable from the true posterior as the numerical error alters the target distribution itself. However, the distortion from the numerical error is unlikely to exhibit consistency over Bayesian updating, and thus posterior SBC would indicate the problem.
@@ 4.3 Posterior SBC in amortized Bayesian inference
In this case study, we show how posterior SBC can serve as a powerful diagnostic for amortized Bayesian inference workflows. Amortized Bayesian inference (see Zammit-Mangion et al.,, 2024, for an overview) constitutes a new wave of Bayesian inference methods, where neural networks learn a direct surrogate for the posterior distribution. Normalizing flows (Papamakarios et al.,, 2021) are arguably the most prominent neural density estimator for amortized inference, but other architectures have recently been explored as well, such as score-based diffusion models (Geffner et al.,, 2023; Sharrock et al.,, 2024), flow matching (Dax et al.,, 2023), and consistency models (Schmitt et al., 2023b, ). For amortized inference, there are no inference diagnostics similar to the convergence diagnostics for MCMC. While posterior SBC for MCMC has the high cost of running MCMC many times, in amortized inference the additional cost of running posterior SBC is negligible.
@@
Amortized inference consists of two stages: (1) A training stage, where a generative neural network $q_{\phi}$ learns to distil relevant information from the probabilistic model based on synthetic data $(\theta,y)\sim p(\theta,y)$ from the joint model; and (2) the inference stage, where the trained neural networks approximate the posterior distribution for an arbitrary new dataset $y_{\mathrm{obs}}$. Amortized inference casts fitting a probabilistic model as a neural network prediction task, which typically takes well under a second. This achieves near-instant approximate posterior sampling $\theta\sim q_{\phi}(\theta\,|\,y_{\mathrm{obs}})$ and even direct posterior density evaluations for arbitrary parameter values $\theta$.
@@
where $\varepsilon\sim\mathcal{N}(0,1)$ is Gaussian noise and $\Delta t$ controls the time resolution (5 milliseconds in our data model). The DDM is additionally parameterized by a decision threshold $\alpha$, an evidence starting point $\beta$ to represent biases, and a non-decision time $\tau$ to account for motor reaction times. We refer to Appendix B in the supplementary material for more details.
In a line of applied work, neurological research has identified neural markers that are associated with decision-making processes. In a recent paper, Ghaderi-Kangavari et al., (2023) propose a set of multiple statistical models that integrate both the cognitive drift-diffusion model and a compressed representation of neural measurements, as found in EEG or fMRI data. Such integrative models, which combine neural processes at the level of single experimental trials, represent the current state-of-the-art in cognitive modeling research. This case study implements two probabilistic models $M_{1}$ and $M_{2}$, which correspond to models #2 and #6 from Ghaderi-Kangavari et al., (2023). We use deep neural networks to fit an amortized posterior approximator for each model. Detailed specifications of models and neural networks can be found in Appendix B in the supplementary material. We showcase how posterior SBC can yield important information to compare the neural approximations of the candidate models conditional on two real datasets from another study (Georgie et al.,, 2018).
@@ Results
In the following, we summarize the main results of the case study and refer to Appendix A in the supplementary material for additional details. In neural amortized inference, we must check the inference validity for different numbers of observations because the latter is passed as a conditioning variable to the inference network $f_{\phi}$. To this end, we confirm that amortized inference on synthetic datasets simulated from the joint model $p(N,\theta,y)$ can recover the ground-truth values of the model parameters that are identifiable for both $N_{\text{obs}}$ and $2N_{\text{obs}}$ (see Appendix A in the supplementary material). Further, we can confirm that prior SBC checking for datasets with $N_{\text{obs}}$ as well as $2N_{\text{obs}}$ yields satisfactory results (see Figure 10). However, this conclusion drastically changes when we perform SBC checking conditional on the real datasets $y_{\text{obs}}^{(1)}$ and $y_{\text{obs}}^{(2)}$. As shown in Figure 10(a), posterior SBC indicates minor data-conditional calibration issues of model $M_{1}$ for the first observed dataset $y_{\text{obs}}^{(1)}$ while calibration conditional on the second observed dataset $y_{\text{obs}}^{(2)}$ is within the acceptance band. The second model $M_{2}$, however, shows pathologically bad calibration of the mean visual encoding latency $\mu_{(e)}$ conditional on either observed dataset (see Figure 10(b)). As a consequence, model $M_{2}$ might be refined according to the iterative Bayesian workflow (Gelman et al.,, 2020; Schad et al.,, 2020) with a particular focus on the mean visual encoding latency parameter $\mu_{(e)}$.
In this case study, posterior SBC could flag calibration issues of an amortized approximator that were not detectable with prior SBC, which illustrates its potential as a default data-conditional diagnostic in amortized Bayesian workflows (Radev et al.,, 2023; Schmitt et al.,, 2024).
@@ 5 Discussion
Prior SBC and posterior SBC have different roles in the Bayesian workflow. Posterior SBC can be used to diagnose whether the used inference algorithm is well calibrated, that is, sampling faithfully from the posterior and the augmented posteriors, which are close to the original posterior. In case of miscalibration, posterior SBC can indicate also the direction and magnitude of the bias for each parameter or other quantity of interest. If there is a worry that the augmented posterior having double data size is too different from the original posterior, it is possible to only use part of the data for the first posterior and make the augmented data size more similar to the original data size.
Passing posterior SBC checking is a necessary but not sufficient condition to trust the posterior inference in the same way as most convergence diagnostics given finite computation time. If the common MCMC diagnostics, like $\widehat{R}$ and divergences,indicate big problems, it is good to first investigate the possible reasons before spending computation time for posterior SBC. Even if the convergence issues are not completely solved, it is possible to use posterior SBC to assess the direction and magnitude of the potential bias.